The following is a straight-line demand curve that confronts a single firm.

 

Quantity

Price demanded (3) (4)   

$6 1 _____ _____

5 2 _____ _____

4 3 _____ _____

3 4 _____ _____

2 5 _____ _____

1 6 _____ _____

 

(a) In column 3, compute total revenue. In column 4, compute the coefficient for the price elasticity of demand at each price using the midpoints formula.


1
Expert's answer
2021-07-25T16:22:55-0400

Solution:

a.). The table is as below:

Total Revenue (TR) = P "\times" Q


Price elasticity of demand (Ped) = "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"